# Zihao Zhang, Guodong Li, and Sihuang Hu

### v1

@@ 3.2 The lower bound on avg-min repair I/O
    Let $a$ be an element in $[t]$. According to Lemma 16 (2), if there exists $b\in[t]$ such that $\langle\mathbf{u}_{b}\rangle=\langle\mathbf{u}_{a}\rangle$, $\langle\mathbf{v}_{b}\rangle\neq\langle\mathbf{v}_{a}\rangle$ or $\langle\mathbf{u}_{b}\rangle\neq\langle\mathbf{u}_{a}\rangle$, $\langle\mathbf{v}_{b}\rangle=\langle\mathbf{v}_{a}\rangle$, then $\overline{\mathcal{N}(\mathbf{F}^{\prime}_{a})}\cap\overline{\mathcal{N}(\mathbf{F}^{\prime}_{b})}\cap[t]=\emptyset$. It follows that $\mathcal{N}(\mathbf{F}^{\prime}_{a})\cup\mathcal{N}(\mathbf{F}^{\prime}_{b})=[t]$, which completes the proof. Otherwise, for any $b\in[t]$, either $\langle\mathbf{u}_{b}\rangle=\langle\mathbf{u}_{a}\rangle$, $\langle\mathbf{v}_{b}\rangle=\langle\mathbf{v}_{a}\rangle$, or $\langle\mathbf{u}_{b}\rangle\neq\langle\mathbf{u}_{a}\rangle$, $\langle\mathbf{v}_{b}\rangle\neq\langle\mathbf{v}_{a}\rangle$. This restricts $(\langle\mathbf{u}_{i}\rangle,\langle\mathbf{v}_{i}\rangle)$ to choosing only two of the four situations. For example, if $(\langle\mathbf{u}_{a}\rangle,\langle\mathbf{v}_{a}\rangle)=(\langle\mathbf{e}_{1}\rangle,\langle\mathbf{e}_{1}\rangle)$, then {(⟨u_i ⟩, ⟨v_i ⟩): i∈[t]}⊆{(⟨e_1 ⟩, ⟨e_1 ⟩),(⟨e_2 ⟩, ⟨e_2 ⟩)} . According to Lemma 16 (1), we have $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|\leq 2$. Recall that in Lemma 14, we have $\bigcup_{i\in[t]}\mathcal{N}(\mathbf{F}^{\prime}_{i})=[t]$. There exist $b\in[t]$ such that $\mathcal{N}(\mathbf{F}^{\prime}_{a})\cup\mathcal{N}(\mathbf{F}^{\prime}_{b})=\bigcup_{i\in[t]}\mathcal{N}(\mathbf{F}^{\prime}_{i})=[t].$
  3. 3.
    Recall that the choices of $(\langle\mathbf{u}_{i}\rangle,\langle\mathbf{v}_{i}\rangle)$ for all $i\in[t]$ are $(\langle\mathbf{e}_{1}\rangle,\langle\mathbf{e}_{1}\rangle)$ $(\langle\mathbf{e}_{1}\rangle,\langle\mathbf{e}_{2}\rangle)$ $(\langle\mathbf{e}_{2}\rangle,\langle\mathbf{e}_{1}\rangle)$ $(\langle\mathbf{e}_{2}\rangle,\langle\mathbf{e}_{2}\rangle)$. By Lemma 16 (1), we have $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|\leq 4$. If $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|=4$, it implies that
    $\{(\langle\mathbf{u}_{i}\rangle,\langle\mathbf{v}_{i}\rangle):i\in[t]\}$
    $=$ $\{(\langle\mathbf{e}_{1}\rangle,\langle\mathbf{e}_{1}\rangle),(\langle\mathbf{e}_{1}\rangle,\langle\mathbf{e}_{2}\rangle),(\langle\mathbf{e}_{2}\rangle,\langle\mathbf{e}_{1}\rangle),(\langle\mathbf{e}_{2}\rangle,\langle\mathbf{e}_{2}\rangle)\}.$
    Assume $\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}=\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4}\}$ and $\mathcal{S}_{1}\prec\mathcal{S}_{2}\prec\mathcal{S}_{3}\prec\mathcal{S}_{4}$. For any $3$ elements of $\{(\langle\mathbf{e}_{1}\rangle,\langle\mathbf{e}_{1}\rangle)$, $(\langle\mathbf{e}_{1}\rangle,\langle\mathbf{e}_{2}\rangle),$ $(\langle\mathbf{e}_{2}\rangle,\langle\mathbf{e}_{1}\rangle),(\langle\mathbf{e}_{2}\rangle,\langle\mathbf{e}_{2}\rangle)\}$, we can find $2$ elements $(\langle\mathbf{e}_{a}\rangle,\langle\mathbf{e}_{b}\rangle),(\langle\mathbf{e}_{c}\rangle,\langle\mathbf{e}_{d}\rangle)$ such that either $\langle\mathbf{e}_{a}\rangle=\langle\mathbf{e}_{c}\rangle$, $\langle\mathbf{e}_{b}\rangle\neq\langle\mathbf{e}_{d}\rangle$ or $\langle\mathbf{e}_{a}\rangle\neq\langle\mathbf{e}_{c}\rangle$, $\langle\mathbf{e}_{b}\rangle=\langle\mathbf{e}_{d}\rangle$, where $a,b,c,d\in[2]$. According to Lemma 16 (2), for any $3$ elements of $\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4}\}$, there must exist a pair of elements whose cup is $[t]$. Thus, we can find two elements of $\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\}$ whose cup is $[t]$. This contradicts Lemma 17 (2). In conclusion, we have $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|\leq 3$.
@@
Recall that
$\bar{\gamma}(\mathcal{C})=\frac{1}{n}{\sum_{i\in[n]}\operatorname{IO}(\mathbf{F}^{\prime}_{i})}=\frac{1}{n}{\sum_{i\in[n]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|}+k,$
where $\mathcal{N}(\mathbf{F}^{\prime}_{i})$ is equal to $\mathcal{N}(\mathbf{F}^{\prime}_{i})$. Since $|\mathcal{N}(\mathbf{F}^{\prime}_{k+1})|\leq|\mathcal{N}(\mathbf{F}^{\prime}_{n})|$ and $\mathcal{N}(\mathbf{F}^{\prime}_{t+1})=\dots=\mathcal{N}(\mathbf{F}^{\prime}_{k+1})$, we have
$\sum_{i\in[n]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|\geq\sum_{i\in[t]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|+(k+2-t)|\mathcal{N}(\mathbf{F}^{\prime}_{k+1})|.$
Lemma 17 (3) tell us that $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|\leq 3$. In the following, we will discuss the value of $\sum_{i\in[t]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|$ case by case.
Case 1: $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|=1$.
It implies that $\mathcal{N}(\mathbf{F}^{\prime}_{1})=\dots=\mathcal{N}(\mathbf{F}^{\prime}_{t})$. Recall that $i\in\mathcal{N}(\mathbf{F}^{\prime}_{i})$, we have $\sum_{i\in[t]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|\geq t^{2}$.
Case 2: $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|=2$.
Assume that $\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}=\{\mathcal{S}_{1},\mathcal{S}_{2}\}$ and $\mathcal{S}_{1}\prec\mathcal{S}_{2}$. Let $\mathcal{T}_{1}=\{i\in[t]:\mathcal{N}(\mathbf{F}^{\prime}_{i})=\mathcal{S}_{1}\}$ and $\mathcal{T}_{2}=\{i\in[t]:\mathcal{N}(\mathbf{F}^{\prime}_{i})=\mathcal{S}_{2}\}$. Recall that $i\in\mathcal{N}(\mathbf{F}^{\prime}_{i})$, we have $\mathcal{T}_{1}\subseteq\mathcal{S}_{1}$, $\mathcal{T}_{2}\subseteq\mathcal{S}_{2}$ and $|\mathcal{T}_{1}|+|\mathcal{T}_{2}|=t$. We can compute that
$\sum_{i\in[t]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|$ $=|\mathcal{T}_{1}||\mathcal{S}_{1}|+|\mathcal{T}_{2}||\mathcal{S}_{2}|\geq|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|^{2}$
$\geq(|\mathcal{T}_{1}|+|\mathcal{T}_{2}|)^{2}/2\text{~~(Cauchy-Schwarz inequality)}$
$\geq t^{2}/2.$
Case 3: $|\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}|=3$.
Assume that $\{\mathcal{N}(\mathbf{F}^{\prime}_{i}):i\in[t]\}=\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3}\}$ and $\mathcal{S}_{1}\prec\mathcal{S}_{2}\prec\mathcal{S}_{3}$. Let $\mathcal{T}_{1}=\{i\in[t]:\mathcal{N}(\mathbf{F}^{\prime}_{i})=\mathcal{S}_{1}\}$, $\mathcal{T}_{2}=\{i\in[t]:\mathcal{N}(\mathbf{F}^{\prime}_{i})=\mathcal{S}_{2}\}$ and $\mathcal{T}_{3}=\{i\in[t]:\mathcal{N}(\mathbf{F}^{\prime}_{i})=\mathcal{S}_{3}\}$. It implies that $\mathcal{T}_{1}\subseteq\mathcal{S}_{1}$, $\mathcal{T}_{2}\subseteq\mathcal{S}_{2}$, $\mathcal{T}_{3}\subseteq\mathcal{S}_{3}$ and $|\mathcal{T}_{1}|+|\mathcal{T}_{2}|+|\mathcal{T}_{3}|=t$. By Lemma 13, we have $(\mathcal{T}_{2}\cup\mathcal{T}_{3})\cap\mathcal{S}_{1}=\emptyset$. And so $\mathcal{S}_{1}=\mathcal{T}_{1}$. According to Lemma 17 (1) and (2), we can get $\mathcal{S}_{1}\cup\mathcal{S}_{3}=[t]$ and $\mathcal{S}_{2}\cup\mathcal{S}_{3}=[t]$. It implies that $t]\setminus\mathcal{S}_{3}\subseteq\mathcal{S}_{1}\cap\mathcal{S}_{2}$. Recall that $\mathcal{T}_{2}\subseteq\mathcal{S}_{2}$, we can get
$|\mathcal{S}_{2}|\geq|([t]\setminus\mathcal{S}_{3})\cup\mathcal{T}_{2}|$ $=|[t]\setminus\mathcal{S}_{3}|+|\mathcal{T}_{2}|-|([t]\setminus\mathcal{S}_{3})\cap\mathcal{T}_{2}|$
$\geq|[t]\setminus\mathcal{S}_{3}|+|\mathcal{T}_{2}|-|\mathcal{S}_{1}\cap\mathcal{T}_{2}|$
$=|[t]\setminus\mathcal{S}_{3}|+|\mathcal{T}_{2}|-|\mathcal{T}_{1}\cap\mathcal{T}_{2}|$
$=|[t]\setminus\mathcal{S}_{3}|+|\mathcal{T}_{2}|.$
Let $c=|[t]\setminus\mathcal{S}_{3}|$. Since $t]\setminus\mathcal{S}_{3}\subseteq\mathcal{S}_{1}\cap\mathcal{S}_{2}$, we have $0\leq c\leq|\mathcal{S}_{1}|=|\mathcal{T}_{1}|$. Note that $|\mathcal{S}_{3}|=t-|[t]\setminus\mathcal{S}_{3}|=t-c$, we can compute that
$~\sum_{i\in[t]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|=|\mathcal{T}_{1}||\mathcal{S}_{1}|+|\mathcal{T}_{2}||\mathcal{S}_{2}|+|\mathcal{T}_{3}||\mathcal{S}_{3}|$
$\geq$ $~|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|(|\mathcal{T}_{2}|+c)+|\mathcal{T}_{3}|(t-c)$
$=$ $~|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|^{2}+|\mathcal{T}_{3}|t+(|\mathcal{T}_{2}|-|\mathcal{T}_{3}|)c$
$\geq$ $~\min\{|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|^{2}+|\mathcal{T}_{3}|t,|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{3}|^{2}+|\mathcal{T}_{2}|t\}$
$\geq$ $~\min\{\frac{(|\mathcal{T}_{1}|+|\mathcal{T}_{2}|)^{2}}{2}+|\mathcal{T}_{3}|t,\frac{(|\mathcal{T}_{1}|+|\mathcal{T}_{3}|)^{2}}{2}+|\mathcal{T}_{2}|t\}$
$=$ $~\min\{\frac{(t-|\mathcal{T}_{3}|)^{2}}{2}+|\mathcal{T}_{3}|t,\frac{(t-|\mathcal{T}_{2}|)^{2}}{2}+|\mathcal{T}_{2}|t\}$
$=$ $~\min\{\frac{t^{2}+|\mathcal{T}_{3}|^{2}}{2},\frac{t^{2}+|\mathcal{T}_{2}|^{2}}{2}\}$
$\geq$ $~\frac{t^{2}+1}{2}.$
From the discussion of Cases 1, 2, and 3, we can know that $\sum_{i\in[t]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|\geq t^{2}/2$. Since $|\mathcal{N}(\mathbf{F}^{\prime}_{k+1})|\geq|\mathcal{N}(\mathbf{F}^{\prime}_{t})|\geq t/2$ and $|\mathcal{N}(\mathbf{F}^{\prime}_{k+1})|\geq k+1-t$, we can compute that
$\sum_{i\in[n]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|$ $\geq\sum_{i\in[t]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|+(k+2-t)|\mathcal{N}(\mathbf{F}^{\prime}_{k+1})|$
$\geq t^{2}/2+(k+2-t)\max\{t/2,k+1-t\}$
$\geq\frac{(k+1)(k+2)}{3}$
Hence, $\bar{\gamma}(\mathcal{C})=\frac{\sum_{i\in[n]}|\mathcal{N}(\mathbf{F}^{\prime}_{i})|}{n}+k\geq\frac{4k+1}{3}$. ∎
4 $(k+2,k,2)$ MDS array codes with optimal repair

### v2

@@ 3.2 The Lower Bound on Avg-Min Repair I/O
    Fix an element $a\in[t]$. We consider two cases for the remaining elements:
    If there exists $b\in[t]$ such that exactly one of the identities $\langle\bm{u}_{b}\rangle=\langle\bm{u}_{a}\rangle$ or $\langle\bm{v}_{b}\rangle=\langle\bm{v}_{a}\rangle$ holds, then by Lemma 16 (2), $\mathcal{N}_{a}\cup\mathcal{N}_{b}=[k]$. This implies $t=k$ as shown in Part 1. Thus, $\mathcal{N}_{a}\cup\mathcal{N}_{b}=[t]$, and the proof is complete.
    Otherwise, for all $b\in[t]$, the pairs $(\langle\bm{u}_{b}\rangle,\langle\bm{v}_{b}\rangle)$ must either match $(\langle\bm{u}_{a}\rangle,\langle\bm{v}_{a}\rangle)$ completely, or differ in both coordinates. This restricts the set of all pairs $\{(\langle\bm{u}_{i}\rangle,\langle\bm{v}_{i}\rangle):i\in[t]\}$ to at most two distinct configurations. By Lemma 16 (1), this implies $|\{\mathcal{N}_{i}:i\in[t]\}|\leq 2$. Since $\bigcup_{i\in[t]}\mathcal{N}_{i}=[t]$, we can always find some $b\in[t]$ such that $\mathcal{N}_{a}\cup\mathcal{N}_{b}=[t]$.
@@
Recall that the average metric is given by
$\bar{\gamma}(\mathcal{C})=\frac{1}{n}\sum_{i\in[n]}\operatorname{IO}(M^{\prime}_{i})=\frac{1}{n}\sum_{i\in[n]}|\mathcal{N}_{i}|+k.$
Given the ordering $|\mathcal{N}_{k+1}|\leq|\mathcal{N}_{n}|$ and the uniform trailing sets $\mathcal{N}_{t+1}=\dots=\mathcal{N}_{k+1}$, we can partition and bound the total sum as:
$\sum_{i\in[n]}|\mathcal{N}_{i}|\geq\sum_{i\in[t]}|\mathcal{N}_{i}|+(k+2-t)|\mathcal{N}_{k+1}|.$
According to Lemma 17, the number of distinct sets within the prefix is strictly bounded by $|\{\mathcal{N}_{i}:i\in[t]\}|\leq 3$. We will analyze the partial sum $\sum_{i\in[t]}|\mathcal{N}_{i}|$ by exhaustively discussing the three possible cases.
Case 1: $|\{\mathcal{N}_{i}:i\in[t]\}|=1$. This implies total uniformity: $\mathcal{N}_{1}=\dots=\mathcal{N}_{t}$. Since every index satisfies $i\in\mathcal{N}_{i}$, the single unique set must contain all $i\in[t]$, meaning its size is at least $t$. Thus,
$\sum_{i\in[t]}|\mathcal{N}_{i}|\geq t\cdot t=t^{2}.$
Case 2: $|\{\mathcal{N}_{i}:i\in[t]\}|=2$. Assume the two distinct sets are $\mathcal{S}_{1}\prec\mathcal{S}_{2}$. We partition the index set $[t]$ into $\mathcal{T}_{1}=\{i\in[t]:\mathcal{N}_{i}=\mathcal{S}_{1}\}$ and $\mathcal{T}_{2}=\{i\in[t]:\mathcal{N}_{i}=\mathcal{S}_{2}\}$. Since $i\in\mathcal{N}_{i}$, it follows naturally that $\mathcal{T}_{1}\subseteq\mathcal{S}_{1}$ and $\mathcal{T}_{2}\subseteq\mathcal{S}_{2}$, with $|\mathcal{T}_{1}|+|\mathcal{T}_{2}|=t$. Applying the Cauchy-Schwarz inequality, we compute:
$\sum_{i\in[t]}|\mathcal{N}_{i}|$ $=|\mathcal{T}_{1}||\mathcal{S}_{1}|+|\mathcal{T}_{2}||\mathcal{S}_{2}|$
$\geq|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|^{2}$
$\geq\frac{(|\mathcal{T}_{1}|+|\mathcal{T}_{2}|)^{2}}{2}=\frac{t^{2}}{2}.$
Case 3: $|\{\mathcal{N}_{i}:i\in[t]\}|=3$. By the property (1) in Lemma 17, having exactly 3 distinct sets forces $t=k$. Let the sets be strictly ordered as $\mathcal{S}_{1}\prec\mathcal{S}_{2}\prec\mathcal{S}_{3}$. We partition the index space $[k]$ into $\mathcal{T}_{j}=\{i\in[k]:\mathcal{N}_{i}=\mathcal{S}_{j}\}$ for $j\in\{1,2,3\}$. This guarantees $\mathcal{T}_{j}\subseteq\mathcal{S}_{j}$ and $\sum_{j=1}^{3}|\mathcal{T}_{j}|=k$.
By Lemma 13, the strictly smaller set $\mathcal{S}_{1}$ cannot contain indices belonging to the larger sets, yielding $(\mathcal{T}_{2}\cup\mathcal{T}_{3})\cap\mathcal{S}_{1}=\emptyset$. Since $\mathcal{T}_{1}\subseteq\mathcal{S}_{1}$ and the partitions form $[k]$, this completely restricts $\mathcal{S}_{1}=\mathcal{T}_{1}$. Furthermore, Lemma 17 dictates the maximal coverages $\mathcal{S}_{1}\cup\mathcal{S}_{3}=[k]$ and $\mathcal{S}_{2}\cup\mathcal{S}_{3}=[k]$, which implies $k]\setminus\mathcal{S}_{3}\subseteq\mathcal{S}_{1}\cap\mathcal{S}_{2}$. Because $\mathcal{S}_{1}=\mathcal{T}_{1}$ and the index partitions are mutually disjoint, we deduce $([k]\setminus\mathcal{S}_{3})\cap\mathcal{T}_{2}\subseteq\mathcal{T}_{1}\cap\mathcal{T}_{2}=\emptyset$. Therefore, utilizing $\mathcal{T}_{2}\subseteq\mathcal{S}_{2}$, we have:
$|\mathcal{S}_{2}|\geq|([k]\setminus\mathcal{S}_{3})\cup\mathcal{T}_{2}|=|[k]\setminus\mathcal{S}_{3}|+|\mathcal{T}_{2}|.$
Let $c=|[k]\setminus\mathcal{S}_{3}|$. The inclusion $k]\setminus\mathcal{S}_{3}\subseteq\mathcal{S}_{1}=\mathcal{T}_{1}$ naturally bounds $c$ within $0\leq c\leq|\mathcal{T}_{1}|$. Noting that $|\mathcal{S}_{3}|=k-c$, we can expand the sum:
$\sum_{i\in[k]}|\mathcal{N}_{i}|$ $=|\mathcal{T}_{1}||\mathcal{S}_{1}|+|\mathcal{T}_{2}||\mathcal{S}_{2}|+|\mathcal{T}_{3}||\mathcal{S}_{3}|$
$\geq|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|(|\mathcal{T}_{2}|+c)+|\mathcal{T}_{3}|(k-c)$
$=|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|^{2}+|\mathcal{T}_{3}|k+(|\mathcal{T}_{2}|-|\mathcal{T}_{3}|)c.$
Notice that the expression is a linear function with respect to $c\in[0,|\mathcal{T}_{1}|]$. Its minimum must therefore occur at one of the boundary points, $c=0$ or $c=|\mathcal{T}_{1}|$. Checking both extremes yields:
$\sum_{i\in[k]}|\mathcal{N}_{i}|$ $\geq\min\Big\{|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{2}|^{2}+|\mathcal{T}_{3}|k,\;|\mathcal{T}_{1}|^{2}+|\mathcal{T}_{3}|^{2}+|\mathcal{T}_{2}|k\Big\}$
$\geq\min\left\{\frac{(|\mathcal{T}_{1}|+|\mathcal{T}_{2}|)^{2}}{2}+|\mathcal{T}_{3}|k,\;\frac{(|\mathcal{T}_{1}|+|\mathcal{T}_{3}|)^{2}}{2}+|\mathcal{T}_{2}|k\right\}.$
By substituting $|\mathcal{T}_{1}|+|\mathcal{T}_{2}|=k-|\mathcal{T}_{3}|$ and $|\mathcal{T}_{1}|+|\mathcal{T}_{3}|=k-|\mathcal{T}_{2}|$, the expression beautifully simplifies to:
$\sum_{i\in[k]}|\mathcal{N}_{i}|$ $\geq\min\left\{\frac{k^{2}+|\mathcal{T}_{3}|^{2}}{2},\;\frac{k^{2}+|\mathcal{T}_{2}|^{2}}{2}\right\}\geq\frac{k^{2}+1}{2}.$
Synthesizing the discussions from Cases 1, 2, and 3, we establish the universal lower bound $\sum_{i\in[t]}|\mathcal{N}_{i}|\geq t^{2}/2$. Given the size properties $|\mathcal{N}_{k+1}|\geq|\mathcal{N}_{t}|\geq t/2$ and $|\mathcal{N}_{k+1}|\geq k+1-t$, we can lower-bound the total sum:
$\sum_{i\in[n]}|\mathcal{N}_{i}|$ $\geq\sum_{i\in[t]}|\mathcal{N}_{i}|+(k+2-t)|\mathcal{N}_{k+1}|$
$\geq\frac{t^{2}}{2}+(k+2-t)\max\left\{\frac{t}{2},k+1-t\right\}$
$\geq\frac{(k+1)(k+2)}{3}.$
Consequently, dividing by $n$ and adding $k$, we obtain the final bound for the average metric:
$\bar{\gamma}(\mathcal{C})=\frac{1}{n}\sum_{i\in[n]}|\mathcal{N}_{i}|+k\geq\frac{4k+1}{3}.$
4 $(k+2,k,2)$ MDS Array Codes with Optimal Pepair